OpenStudy (anonymous):

Events A and B are mutually exclusive. Also, P(A) = 0.12 and P(B) = 0.25. What is P(A or B)? A. 0.03 B. 0.25 C. 0.13 D. 0.37

OpenStudy (ilovecake):

D.0.37

OpenStudy (amorfide):

P(AUB)= P(A)+P(B)-P(AnB)

OpenStudy (ilovecake):

If A and B are mutually exclusive, then p(A and B) = 0. So p(A or B) = p(A) + p(B) - p(A and B) = 0.12 + 0.25 - 0 = 0.37 Hope that helps.

OpenStudy (anonymous):

it does, thanks